woodchuck runs 20 m to the right in 5 s, then turns and runs 10 m to the left in 3 s. (a) What is the average velocity of the woodchuck? (b) What is its average speed?

1
Expert's answer
2022-03-03T12:05:16-0500

(a) The average velocity of the woodchuck

"v_{ave}=\frac{d_1+d_2}{t_1+t_2}=\frac{20-10}{5+3}=1.25\:\rm m/s\; to \:the\: right"

(b) The average speed of the woodchuck

"v_{ave}=\frac{|d_1|+|d_2|}{t_1+t_2}=\frac{20+10}{5+3}=3.75\:\rm m/s"